### v2

# Fluid dynamics, Statistical physics, Theoretical physics

Abstract
Strongly nonlinear dynamics, from fluid turbulence to quantum chromodynamics, have long constituted some of the most challenging problems in theoretical physics. This review describes a unified theoretical framework, the loop space calculus, which offers an analytical approach to these problems. The central idea is a shift in perspective from pointwise fields to integrated loop observables, a transformation that recasts the governing nonlinear equations into a universal linear diffusion equation in the space of loops. This framework, supported by recent mathematical analysis, is analytically solvable and yields an exact, parameter-free solution for decaying hydrodynamic turbulence—the Euler ensemble—which is shown to be dual to a solvable string theory. The theory’s predictions include: (i) the unification of spatial and temporal scaling laws, governed by two related, infinite spectra of intermittency and decay exponents derived from the nontrivial zeros of the Riemann zeta function; (ii) a first-order phase transition in magnetohydrodynamic (MHD) turbulence; and (iii) the formation of quantized, concentric shells in passive scalar mixing. The theory also predicts log-periodic oscillations in correlation functions — effects not captured by standard phenomenology — for which there is now emerging evidence from high-precision turbulence experiments. The appearance of identical mathematical structures as solutions to the turbulent regime of Yang–Mills gradient flow points to the broad applicability of this approach. The framework also yields a new type of analytic Hodge-dual matrix surface that solves the Yang–Mills fixed-point loop equation by harmonic map, opening the way for a geometric formulation of QCD string theory.
keywords:
turbulence, Navier-Stokes, MHD, Yang-Mills theory, loop equations, string theory duality, Riemann zeta function
1 Introduction
In his 1964 Lectures [Feynman et al.(2011)Feynman, Leighton, and Sands], Richard Feynman identified the analysis of turbulent fluids as a problem "left over from a long time ago… Nobody in physics has really been able to analyze it mathematically satisfactorily." Decades later, this challenge, along with the related problem of non-perturbative dynamics in Yang-Mills gauge theory, has remained a frontier of theoretical science. These phenomena are governed by strongly nonlinear equations that defy standard analytical techniques, forcing a reliance on phenomenological models or numerical simulation.
An analytical approach to these problems emerges from a new theoretical paradigm, which is the subject of this review. The core idea is to shift the descriptive language of physics away from local, pointwise fields (like velocity $\boldsymbol{v}(\boldsymbol{r})$ or the gauge potential $A_{\mu}(x)$) and towards non-local, integrated observables defined on closed loops (like velocity circulation $\Gamma_{C}$ or the Wilson loop $W[C]$).
This reformulation is not merely a change of variables; it constitutes a dimensional reduction of the dynamics that also eliminates the nonlinearity. It transforms the governing partial differential equations into a universal, linear diffusion equation in the space of loops, rendering the statistical evolution analytically tractable via a functional Fourier transformation to momentum loop space, where the loop-space diffusion reduces to an algebraic problem. This evolution converges to a universal attractor — the Euler ensemble — which is dual to a solvable string theory.
A central result of this framework is its prediction of two distinct but related spectra of exponents. For the spatial scaling of velocity correlations, where phenomenological models typically assume a single dominant power law for each statistical moment, the theory reveals an infinite, discrete spectrum of intermittency exponents. The leading exponent in this series, which governs the dominant large-scale behavior, can be regarded as the "ground state" of the scaling law. The subsequent exponents then define a hierarchy of corrections to this leading behavior, akin to quantum corrections, which become relevant at finer scales. In the same manner, the theory determines a full spectrum of decay exponents for the temporal evolution of integrated quantities, such as the kinetic energy. Here too, the principal exponent, which corresponds to the observed decay law ($E\propto t^{-5/4}$), acts as the ground state, with the rest of the spectrum describing higher-order corrections to the temporal decay. Crucially, these two spectra, traditionally considered independent, emerge from the same underlying number-theoretic structure of the Euler ensemble. This solution establishes a direct connection between the fine-scale spatial statistics and the long-term temporal decay of the turbulent flow.
The Euler ensemble solution provides an explicit analytic expression for the full spectrum of these exponents, relating them to rational numbers and the complex zeros of the Riemann zeta function. The imaginary parts of the complex exponents predict oscillations in correlation functions on a logarithmic scale, a theoretical result consistent with recent observations from wind-tunnel experiments by the Max Planck turbulence group (Fig. 6).
A Brief Historical Context
The challenge of deriving a statistical theory of turbulence directly from the Navier-Stokes equations has a long history. Notably, Eberhard Hopf pioneered the use of functional methods, introducing an equation for the characteristic functional of the velocity field and envisioning a universal turbulent attractor governing the long-time statistics (see a recent review [Ohkitani(2020)]). However, Hopf’s functional equation proved intractable, leaving the nature of the attractor unclear for decades. Now, with the loop equation replacing the Hopf functional equation, we have its solution —the Euler ensemble —as the realization of Hopf’s conjecture about the turbulent attractor.
Later, the loop space formalism emerged, rooted in the Makeenko-Migdal (MM) equations of the late 1970s and early 1980s [Makeenko and Migdal(1979), Makeenko and Migdal(1981), Migdal(1983)]. These were immediately recognized in the high-energy physics community as the first exact, non-perturbative dynamical equations for QCD. However, the initial hope of finding an analytical solution for quark confinement soon faded. The mathematical formulation, while geometrically insightful, proved analytically intractable, hampering further progress and shifting the community’s focus toward lattice gauge theory. The loop approach offered a different path compared to Hopf’s formulation, focusing on circulation observables.
The scope of the formalism was extended to fluid dynamics in the early nineties with the derivation of the loop equations for the Navier-Stokes system [Migdal(1995b)]. Within this framework, a WKB solution was derived for the limit of large, smooth loops. This solution established a theoretical "area law" for the statistics of velocity circulation. A transformation to a dual "momentum loop space" was also shown to reduce the problem to a singular, one-dimensional equation. For decades, however, further progress was impeded by the mathematical challenges posed by this equation and by a lack of high-precision data with which to test the theory’s predictions.
The theoretical landscape evolved considerably after 2019, when high-resolution direct numerical simulations (DNS) by Sreenivasan and collaborators reported results consistent with the area law derived nearly three decades earlier [Iyer et al.(2021)Iyer, Bharadwaj, and Sreenivasan, Pumir(2021)]. This numerical evidence renewed interest in the approach and provided the impetus to resolve the long-standing mathematical challenges of the momentum loop equation. The complete analytical solution that emerged, realizing aspects of Hopf’s original vision via loop space, is the subject of this review. (See also [Migdal(2024h)] for further discussion on the connection to Hopf’s work).
This history underscores a critical point: a correct, nonsingular, and tractable loop space formalism was the missing key. This review introduces this key—a new, nonsingular calculus that renders the loop equations solvable.
Recent Mathematical and Numerical Validation
This theoretical structure, initially derived from physical principles, has recently received significant independent validation. From the number theory perspective, the statistical properties of the Euler ensemble at finite $N$ were studied in [Basak and Zaharescu(2025)] by combinatorial methods, confirming and extending the results and conjectures made in the first paper [Migdal(2023)]. In addition, the distribution of radii of regular star polygons with unit side lengths was recently reproduced by Debmalya Basak using rigorous methods of number theory (private communication, to be published). From a Navier-Stokes dynamics perspective, a rigorous analysis by Brue and De Lellis [Elia and Lellis(2025)] has confirmed that the Euler ensemble is an exact solution to the discrete version of the loop equation (Theorem 10.6 in their paper). While the final step—proving the existence of the continuum limit of the fluid dynamics observables as the number of loop segments $N\to\infty$—remains an open mathematical problem, their result provides a firm mathematical foundation for the core of the theory. It confirms that the Euler ensemble is not an approximation but an exact solution at the level of the regularized theory.
This mathematical validation is now strongly complemented by new high-resolution 4K DNS results from Sreenivasan and collaborators [Sreenivasan and Rodhiya, Akash(2026)]. These results verify the theory’s quantitative predictions with high precision, representing a significant development. As discussed in Sections 4.1 and 8, the DNS data for key observables—such as the $E\propto t^{-5/4}$ decay law and the universal, nonlinear effective index—are in excellent, parameter-free agreement with the theoretical curves derived from the Euler ensemble (see Fig. 2). Together, these mathematical and numerical results provide strong support for the loop space calculus in the Navier-Stokes equation, confirming the Euler ensemble as a robust solution for decaying turbulence.
Outline of the Review
This review is organized as follows. The discussion begins with an introduction to the general loop space approach, applicable to both abelian fields in fluid dynamics and nonabelian fields in Yang-Mills theory. The approach is then applied to a hierarchy of turbulence problems, beginning with the foundational case of decaying hydrodynamic turbulence, then the transport of passive scalars, and finally the more complex case of magnetohydrodynamic (MHD) turbulence. A subsequent section situates these analytical solutions within the field’s historical context and clarifies their relationship to classical paradigms of turbulence theory. A particularly striking consequence of this framework emerges in gauge theory. In the Yang–Mills gradient flow, the same loop-space formalism leads to an exact analytic solution—the Hodge-dual matrix surface—which solves the fixed-point Yang–Mills loop equation by harmonic map [Migdal(2025a)]. This result establishes a concrete geometric foundation for the QCD confining string and provides a direct analytical bridge between turbulence, gauge theory, and string dualities.
2 The Loop Space Approach: From Nonlinear PDEs to Linear Diffusion
The foundation of our approach is the loop functional, $\Psi[C]$, which acts as‘ the characteristic function for the probability distribution of a loop observable. For fluid dynamics, this is the circulation, $\Gamma_{C}=\oint_{C}\boldsymbol{v}\cdot d\boldsymbol{r}$, and the functional is $\Psi[C]=\langle\exp\left(\mathrm{i}\Gamma_{C}/\nu\right)\rangle$. For gauge theory, it is the Wilson loop.
Navier-Stokes circulation equation
Let us start with the Navier-Stokes equation for an incompressible velocity field
$\partial_{t}v_{\alpha}=-v_{\beta}\partial_{\beta}v_{\alpha}+\nu\partial_{\beta}\partial_{\beta}v_{\alpha}-\partial_{\alpha}p;\;\partial_{\beta}v_{\beta}=0;$ (1)
This equation leads to the well-known expression for the time derivative of the circulation, which we write in geometric form using the covariant derivative operator
$D_{\mu}=\partial_{\mu}+\frac{\mathrm{i}v_{\mu}}{\nu};$ (2)
Here is the covariant equation for the time derivative of the circulation
$\partial_{t}\Gamma=\nu\oint_{C}dx_{\alpha}\left(\left[D_{\beta},\omega_{\beta\alpha}\right]+v_{\beta}\omega_{\beta\alpha}\right);$ (3)
$\omega_{\mu\nu}=\partial_{\mu}v_{\nu}-\partial_{\nu}v_{\mu};$ (4)
The covariant derivative operator $D_{\mu}$ will play an important role in the loop dynamics. Note that the imaginary unit in front of the velocity field is similar to that in the Abelian gauge theory covariant derivative, keeping the covariant derivative operator anti-Hermitian $D_{\mu}^{\dagger}=-D_{\mu}$. The vorticity itself can be represented as a commutator.
$\mathrm{i}\omega_{\mu\nu}=\nu[D_{\mu},D_{\nu}];$ (5)
The last term in (3) can be eliminated by switching to the "liquid loop" - the one with each point moving with the local velocity. For such a loop, the Kelvin theorem states that the advection term $v_{\beta}\omega_{\beta\alpha}$ is exactly canceled by the term coming from the loop motion.
$\partial_{t}C_{\alpha}(\theta)=v_{\alpha}(C(\theta),t)$ (6)
As a result, the circulation $\tilde{\Gamma}$ of the moving loop satisfies a simple geometric equation
$\mathrm{i}\partial_{t}\tilde{\Gamma}=\nu^{2}\oint_{C}dx_{\alpha}\left[D_{\beta},[D_{\beta},D_{\alpha}]\right]$ (7)
after which the Navier-Stokes equation for liquid loop functional $\tilde{\Psi}$ takes a purely geometric form
$\partial_{t}\tilde{\Psi}[C]=\nu\oint_{C}dx_{\alpha}\left[D_{\beta},[D_{\beta},D_{\alpha}]\right]\tilde{\Psi}[C];$ (8)
Yang-Mills gradient flow
Here is the definition of the Yang-Mills gradient flow
$\partial_{\tau}A_{\nu}=\alpha[D_{\mu},F_{\mu\nu}],$ (9)
$F_{\mu\nu}=[D_{\mu},D_{\nu}];$ (10)
$D_{\mu}=\partial_{\mu}+A_{\mu};$ (11)
Here, the gauge field $A_{\mu}$ belongs to some Lie algebra, $A_{\mu}=\mathrm{i}\sum T_{a}A^{a}_{\mu}$, which we do not need to specify. The equation for the trace of the ordered path integral
$W[C]=\hbox{tr}\,\mathbb{P}\exp\left(\int_{C}dx_{\alpha}A_{\alpha}\right)$ (12)
has a similar geometric form
$\partial_{\tau}W[C]=\alpha\hbox{tr}\,\mathbb{P}\int_{C}dx_{\alpha}\left[D_{\beta},[D_{\beta},D_{\alpha}]\right]\exp\left(\int_{C}dx_{\alpha}A_{\alpha}\right)$ (13)
This formula does not represent a closed equation for the loop functional, as the right-hand side still depends on the dynamic variables in coordinate space, satisfying the nonlinear partial differential equations we intended to solve in the first place.
Operator identity transforms nonlinear flow into loop space diffusion
The transformation of both equations to the loop space diffusion equation is based on the operator calculus invented by Feynman [Feynman(1951)]. As applied to our case of the loop functionals for fluid dynamics and the gauge theory, it allows us to write both of them as the path ordered exponential times the unit operator
$\tilde{\Psi}[C]\otimes\mathbb{I}=\mathbb{P}\exp\left(\oint dx_{\alpha}D_{\alpha}(x_{0})\right);$ (14)
$W[C]\otimes\mathbb{I}=\hbox{tr}\,\mathbb{P}\exp\left(\oint dx_{\alpha}D_{\alpha}(x_{0})\right);$ (15)
This fundamental geometric identity, proven in the [Migdal(2025b)] and our Appendix A, is identical in abelian and non-abelian cases, except that a group space trace is taken in the Yang-Mills case. In both cases, it reveals the geometric meaning of the loop functional as a parallel transport of the covariant derivative operator around a loop $C$, reducing to a c-number for a closed loop. The reduction of the ordered exponential to a c-number is also supported by the Magnus expansion [Magnus(1954)]. This distinction is subtle but critical. In much of the mathematical literature such as [Göckeler and Schücker(1987)], the Wilson loop (the L.H.S. of our identity) is taken as the definition of parallel transport. Our proof, however, establishes this correspondence as a rigorous equality derived from the more fundamental, anchored operator (the R.H.S.). This approach mirrors a historical pattern in theoretical physics where a concept, once viewed as a powerful analogy, is later shown to be a provable, first-principles identity.
Now, the covariant derivative operator $D_{\mu}$ in both theories can be replaced by a functional derivative acting on the ordered exponential involving that same operator (remember that the operators $D_{\mu}(x_{0})$ in this identity refer to the (arbitrary) origin $x_{0}=C(\theta=0)$ on the loop. In this operator form, the functional derivatives $\frac{\delta}{\delta\dot{C}_{\mu}(t\pm)}$ bring down the covariant derivative operator before or after the ordered product, which makes no difference due to the cyclic symmetry of the trace:
$\frac{\delta}{\delta\dot{C}_{\mu}(t\pm)}W(C(.),\tau)=\left\langle\frac{1}{N}\hbox{tr}\,D_{\mu}(C(t\pm))\right$.
$\left.\mathbb{P}\exp\left(\int_{t}^{t+2\pi}ds\dot{C}_{\mu}(s)D_{\mu}(x_{0})\right)\right\rangle$ (16)
In this formula, we used cyclic symmetry of the trace and invariance of the Wilson loop with respect to the choice of the origin on a circle $(0,2\pi)\Rightarrow(t,t+2\pi)$. We call such functional derivatives the dot derivatives.
Applying the same formula for the dot derivative three times and properly contracting tensor indices, we bring the triple commutator $\left[D_{\beta},[D_{\beta},D_{\gamma}]\right]$ from the exponential, which results in the diffusion equation for $\tilde{\Psi}[C]$:
$\partial_{t}\tilde{\Psi}[C,t]=\nu\mathcal{L}_{C}\tilde{\Psi}[C,t]$ (17)
with the diffusion operator
$\mathcal{L}_{C}=\oint d\theta\dot{C}_{\nu}(\theta)\hat{L}_{\nu}(\theta);$ (18)
$\hat{L}_{\nu}(\theta)=T^{\alpha\beta\gamma}_{\nu}\frac{\delta^{3}}{\delta\dot{C}_{\alpha}(\theta-0)\delta\dot{C}_{\beta}(\theta)\delta\dot{C}_{\gamma}(\theta+0)};$ (19)
$T^{\alpha\beta\gamma}_{\nu}=\delta_{\alpha\beta}\delta_{\gamma\nu}+\delta_{\gamma\beta}\delta_{\alpha\nu}-2\delta_{\alpha\gamma}\delta_{\beta\nu};$ (20)
All three arguments of the third functional derivative tend to $\theta$ in the specified order. The dot derivatives generate covariant derivatives inside the ordered product in the Wilson loop, and contraction with the tensor $T$ arranges these operators into a triple commutator
$T^{\alpha\beta\gamma}_{\nu}D_{\alpha}D_{\beta}D_{\gamma}=\left[D_{\mu},\left[D_{\mu},D_{\nu}\right]\right];$ (21)
A Note on Dot Derivatives and Kinematical Discontinuities
It is important to clarify the nature of the "dot derivatives" ($\delta/\delta\dot{C}(\theta)$) that form the basis of loop space calculus. These are the standard functional derivatives with respect to the velocity of a periodic trajectory, familiar from the variational principles of classical mechanics (e.g., the Euler-Lagrange equations).
The profound subtlety arises when multiple derivatives are taken at the same parameter point $\theta$ along the loop. The result depends on the order in which the limit of coincident points is taken. For instance, the action of a second derivative depends on whether we take the points as $(\theta,\theta+\epsilon)$ or $(\theta+\epsilon,\theta)$ in the limit $\epsilon\to 0$.
This ordering dependence results in a finite discontinuity. This discontinuity is not a pathology or a short-distance singularity of the underlying field that needs to be regularized. Rather, it is a fundamental and universal kinematical property of loop functionals.
The difference between the two possible orderings of a second dot derivative is directly proportional to the commutator of the underlying covariant derivatives. This mathematical structure precisely encodes the essential physics:
  • In fluid dynamics, this commutator is the vorticity tensor, $\omega_{\mu\nu}$.
  • In Yang-Mills theory, it is the field strength tensor, $F_{\mu\nu}$.
Thus, the non-commutativity of the dot derivatives at a single point is the loop-space representation of the local field strength. This effect would be present even for the simplest non-trivial fields, such as a constant field strength in Yang-Mills theory or a rigid body rotation ($\boldsymbol{v}=\frac{1}{2}\boldsymbol{\omega}\times\boldsymbol{r}$) in hydrodynamics, where the fields themselves are perfectly smooth. These discontinuities are the mechanism by which the local, rotational dynamics of the field are captured by the calculus on the one-dimensional loop.
Another subtlety is the operator nature of these dot derivatives. Every functional derivative $\delta/\delta\dot{C}(\theta)$ brings down from the path-ordered exponential a covariant derivative operator. The resulting object is no longer a c-number, as this covariant derivative acts on all the factors to the right in the path-ordered exponential. After commuting with all these factors in the ordered product $(1+d\theta\dot{C}_{\mu}(\theta)D_{\mu}(x_{0}))$, this covariant derivative operator moves to the end of the path ordered product and "hangs" there, making the whole expression an operator in Hilbert space rather than a c-number.
The c-number is restored after taking the discontinuity of the first dot derivative or the antisymmetrization by tensor indices of the second dot derivative. After this antisymmetrization, the product of two operators $D_{\mu}D_{\nu}$ in front of the path-ordered exponential becomes a c-number commutator $D_{\mu},D_{\nu}]$, and it no longer differentiates the path-ordered exponential to its right. Now we get back the original Wilson loop and its dot derivatives.
So, we hop into a Hilbert space at each dot derivative but come back to a world of numbers after taking discontinuity or antisymmetrization. This transformation also applies to the triple commutator: it is a c-number only after the symmetrization over tensor indices by multiplication by the tensor $T^{\alpha\beta\gamma}_{\nu}$.
The universal loop equation with varying initial data
This equation governs the evolution of the system’s complete statistical state.
With finite viscosity, there is a universal diffusion equation in loop space. The turbulent limit $\nu\to 0$ corresponds to the WKB limit in that diffusion equation, and it is determined by the zero modes of the operator $\mathcal{L}_{C}$.
In the case of gauge theory, the same equation describes the evolution of the Wilson loop in Yang-Mills gradient flow. The difference lies in the initial data. The general solution of the loop equation in both cases can be written as
$\Psi[C]=\exp\left(\nu t\mathcal{L}_{C}\right)\Psi_{0}[C];$ (22)
$W[C]=\exp\left(\alpha t\mathcal{L}_{C}\right)W_{0}[C]$ (23)
The initial data are, of course, different, corresponding to the distributions of the two different vector fields.
$\Psi_{0}[C]=\left\langle\exp\left(\frac{\mathrm{i}\oint_{C}dx_{\mu}v^{0}_{\mu}(x)}{\nu}\right)\right\rangle_{v^{0}};$ (24)
$W_{0}[C]=\left\langle\hbox{tr}\,\mathbb{P}\exp\left(\oint_{C}dx_{\mu}A^{0}_{\mu}(x)\right)\right\rangle_{A^{0}}$ (25)
The mathematical tool that enables this formulation is a loop space calculus. Unlike previous approaches that required adding singular cusps to loops, our calculus operates entirely within the manifold of smooth loops. Variational operators are defined via derivatives with respect to the loop’s velocity profile, $\dot{C}(\theta)$. This method is free of the singularities and ambiguities that plagued earlier attempts, yielding a well-defined operator $\mathcal{L}_{C}$. The technical details of this calculus are provided in Appendix A.
The most important aspect of this new loop equation is that it is analytically solvable by functional Fourier transform in loop space.
3 Momentum loop equation
Now let us return to the nonsingular loop equation (17). The diffusion equation in any linear space can be exactly solved by the Fourier transform, and the loop space is no exception. However, its infinite dimensionality makes the solution quite nontrivial.
Let us write down an Ansatz for a solution in the form of the momentum loop equation (MLE)
$W[C,\tau]=\left\langle\exp\left(\mathrm{i}\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)P_{\mu}(\theta,\tau)\right)\right\rangle_{P(\tau)};$ (26)
Substituting this Anzatz into the loop equation (17), we find that this equation requires the following evolution of momentum loop
$\frac{\partial_{\tau}\bar{P}}{\nu}=(\Delta P\cdot\bar{P})\Delta P-(\Delta P)^{2}\bar{P};$ (27)
$\bar{P}=\frac{P(\theta+)+P(\theta-)}{2};$ (28)
$\Delta P=P(\theta+)-P(\theta-);$ (29)
The brackets $\left\langle\right\rangle_{P(\tau)}$ correspond to the averaging over an ensemble of solutions of the time evolution of $P_{\mu}(.,\tau)$ described by the above equation.
This Anzatz (26) is the loop space version of the plane wave. The loop equation (17) involves only dot functional derivatives; therefore, an Anzatz would exactly satisfy the loop equation, with each dot derivative $\frac{\delta}{\delta\dot{C}_{\nu}(\theta)}$ equivalent to multiplication of $\mathrm{i}P(\theta)$. There is a one-to-one correspondence between the original Navier-Stokes equation and this algebraic relation between momentum loops and their discontinuities. The nonlinearity of the Navier-Stokes equation results in nonlinearity in the momentum loop equation, with the different terms in the Navier-Stokes equation combined into a simple cubic polynomial, reflecting the structure of the triple commutator of covariant derivatives.
The relation between the loop operator and the discontinuity of the momentum loop $\Delta P_{\nu}(\theta)$ was discovered and investigated in earlier papers in QCD [Migdal(1986), Migdal(1995a), Migdal(1998)]. The Momentum loop equation for QCD involves contact terms that provide boundary conditions in loop space for self-intersecting loops. We are not studying these terms in the present paper, restricting ourselves to the Yang-Mills gradient flow. For QCD, this equation corresponds to the WKB limit of large loops without self-intersections, when some form of the area law is expected as an asymptotic solution.
4 Application I: Decaying Hydrodynamic Turbulence
We first apply the framework to decaying homogeneous isotropic turbulence—Feynman’s "oldest unsolved problem." The momentum loop equation (27) admits a universal, long-time attractor solution that is independent of initial conditions. The time dependence immediately follows from the fact that the right side of (27) is a cubic homogeneous functional of $P$, and the left side is the time derivative of $P$. The equation then becomes a nonlinear relation for the coefficient in front of the power factor
$P_{\mu}(\theta,\tau)=\frac{f_{\mu}(\theta)}{\sqrt{2\nu(\tau+\tau_{0})}};$ (30)
$\bar{f}\left(\Delta f^{2}-1\right)=(\Delta f\cdot\bar{f})\Delta f$ (31)
For the loop functional, it reduces to the following.
$\tilde{\Psi}[C]=\left\langle\exp\left(\frac{\mathrm{i}}{\sqrt{2\nu(\tau+\tau_{0})}}\sum_{k}f_{\mu}(\theta_{k})\Delta C_{\mu}(\theta_{k})\right)\right\rangle_{\mathcal{E}}$ (32)
where the averaging is over the solutions of the recurrent equation (31). The equally spaced points on a unit circle $\theta_{k}=\frac{2\pi k}{N}$ approximate the infinitesimal step $\theta\pm 0$ involved in the original loop equation. Such a function $f(\theta)$, with discontinuity at every angle, can exist only in the sense of distributions, which is why this limiting procedure is required. These subtle issues of discretization of the loop equation are discussed in the recent paper [Elia and Lellis(2025)], where they are justified under certain assumptions about the continuity of the loop and the flow.
The limit $N\to\infty$ should be taken in the end at fixed turbulent viscosity $\tilde{\nu}=\nu N^{2}$. The last relation was derived in [Migdal(2023)] and used later in [Migdal(2024h)]. It guarantees a finite limit for observable correlation functions, which was analytically computed in these papers, using QFT methods combined with number theory.
Recurrent equation on a circle
The last equation (31) for $f_{\mu}(\theta)$ can be solved exactly, by a certain limiting procedure. First, we observe that the left and right sides are vector products with some scalar coefficients. Unless both of these coefficients vanish, these vectors are collinear. But in that case, the area derivative of the Wilson loop identically vanishes at every point on an arbitrary loop $C$. Such would be a trivial solution in which the gauge field is a pure gauge (a velocity field is purely potential, without any vorticity that could lead to turbulence).
The nontrivial solutions, with finite vorticity, all correspond to both scalar coefficients vanishing for every angle $\theta$. We can rewrite these equations as
$(f(\theta+)+f(\theta-))\cdot(f(\theta+)-f(\theta-))=0;$ (33)
$(f(\theta+)-f(\theta-))^{2}=1;$ (34)
The first equation can be rewritten as the continuity of the length of the vector $f(\theta)$
$f(\theta+)^{2}=f(\theta-)^{2}$ (35)
We conclude that these vectors are located on a sphere:
[…]
The second equation relates the radius of the sphere to the constant angle between the consecutive vectors
$1-n(\theta-)\cdot n(\theta+)=\frac{1}{2R^{2}}=\textit{ const }{}$ (37)
There is also an important requirement of periodicity
[…]
This solution in three dimensions is the Euler ensemble.
The Euler Ensemble: A Universal Turbulent Attractor
In this solution, $\boldsymbol{f}(\theta)$ is a universal fractal curve constructed as the continuum limit ($N\to\infty$) of a random walk on a regular star polygon $\{q/p\}$. Its vertices are given by:
$\boldsymbol{f}\left(\frac{2\pi k}{N}\right)=\hat{\Omega}\cdot\{R\cos(\alpha_{k}),R\sin(\alpha_{k}),0\};$ (39)
$R=\frac{1}{2\sin(\beta/2)};$ (40)
where the rotation matrix $\hat{\Omega}\in SO(3)$, the angle step $\beta=2\pi p/q$ with $p/q\in\mathbb{Q}$, and the cumulative angle $\alpha_{k}=\beta\sum_{l=1}^{k}\sigma_{l}$ with $\sigma_{l}=\pm 1$ define the random walk.
The recurrent equations (33) are satisfied in an obvious geometric way: with this radius $R$ the sides have unit length:
$R\left|e^{\mathrm{i}\alpha_{k+1}}-e^{\mathrm{i}\alpha_{k}}\right|=R\left|e^{\mathrm{i}\beta\sigma_{k}}-1\right|=1.$
The Euler ensemble is not a single configuration but a degenerate manifold of solutions, a "fixed stochastic trajectory." The statistical state of turbulence is a distribution over all choices of $(\hat{\Omega},p,q,\{\sigma_{k}\})$.
The periodicity condition $\alpha_{N}-\alpha_{0}=\beta\sum\sigma_{k}=2\pi r$ imposes a number-theoretic constraint on the geometry of the walk: for a walk to close after traversing N steps, the rotation angle $\beta$ must be a rational multiple of $2\pi$, i.e., $\beta=2\pi p/q$. The requirement that the polygon $\{q/p\}$ be irreducible (i.e., not a simpler polygon traced multiple times) restricts $p$ and $q$ to be coprime, meaning the number of unique star polygons for a given denominator $q$ is precisely $\varphi(q)$, the Euler totient function.
This point is crucial– the origin of the link between turbulence and number theory. We are limited to the manifold $\mathbb{Q}$ of rational numbers, which eventually leads to all the "quantum" effects in the solution of decaying turbulence. The statistics of turbulence are thus described by a uniform distribution over all such polygons and all possible random walks on them. This discrete, number-theoretic structure is the hallmark of spontaneous quantization.
The distribution of the variable $X(p,q)=\cot(\pi p/q)^{2}/N^{2}=\left(4R^{2}-1\right)/N^{2}$ for co-prime $1\leq p<q<N$ was computed in [Migdal(2024h)] by advanced number theory methods. In the statistical limit $N\to\infty$, this distribution remains discontinuous
$f(X)=\left(1-\frac{\pi^{2}}{675\zeta(5)}\right)\delta(X)+\frac{\pi^{3}}{3}X\sqrt{X}\Phi\left(\left\lfloor\frac{1}{\pi\sqrt{X}}\right\rfloor\right);$ (41)
with $\Phi$ the totient summatory function:
$\Phi(q)=\sum_{n=1}^{q}\varphi(n);$ (42)
$\varphi(m)=m\prod_{p|m}\left(1-\frac{1}{p}\right)$ (43)
The discontinuities of $f(X)$ at quantized values are related to famous Euler totients $\varphi(n)$
$\Delta f(X_{n})=\frac{\varphi(n)}{3n^{3}};$ (44)
$X_{n}=\frac{1}{\pi^{2}n^{2}}$ (45)
This distribution has finite support $0<X<\frac{1}{\pi^{2}}$ and condenses to a power law as $f(X)\propto\sqrt{X}$ at $X\to 0$.
The Euler ensemble is remarkably universal. As we show in the Appendix, it also describes the solution of the loop equation for the static loop (not moving with the flow). Explicit computation shows that the advection term $v\omega$ present in the static loop equation, when integrated over the loop using the solution for vorticity and (by inverting the curl operator) for velocity for the Euler ensemble, becomes a total derivative of a periodic function, so that the integral over $\theta$ yields zero. This nontrivial cancellation was not built into the solution for the liquid loop but emerged as a free bonus, suggesting a deeper reason for this solution – not yet completely understood.
Historical note (Wikisource)
The study of star polygons dates to antiquity with Pythagoras and his school. The subject was later notably taken up by Thomas Bradwardine, whose 14th-century mathematical work on these figures, conducted while he was Archbishop of Canterbury, is a testament to the era’s polymathic tradition. The statistical limit of random walk on regular star polygons has never been solved before, to the best of our knowledge.
Figure 1: Examples of the regular star polygons, $\{q/p\}$, that form the discrete target space of the dual string theory. The turbulent state, or Euler ensemble, is a statistical average over random walks on all such polygons.
Figure 2: Verification of the Euler-ensemble theory against new high-resolution 4K DNS data [Sreenivasan and Rodhiya, Akash(2026)]. (Top Left): The predicted parabolic relation between time $t$ and the integral length scale $L_{M}(t)$. The simulation data is in excellent agreement with a parabolic fit. (Top Right): The parameter-free prediction for the effective index $f(x)=\langle r\partial_{r}\log(\Delta v^{2})\rangle$. The theory captures the universal curvature seen in the DNS data. (Bottom): Verification of the full energy decay law at $Re_{\lambda}=145$. The DNS data (blue line) deviates from the leading-order term of the theory ($E\propto L_{M}^{-5/2}$, gray dash-dotted line). However, it is in excellent agreement with the complete theoretical solution (orange line), which includes all sub-leading exponents from the Mellin-Barnes integral. Only the two unknown scales (energy and length) were fitted to match the DNS data with the theory; no dimensionless parameters were changed. Such a perfect fit with only two parameters for about a thousand data points provides a high level of confidence.
4.1 Key Prediction 1: The Energy Spectrum and Intermittency
The Euler ensemble provides a complete, parameter-free prediction for all statistical properties of decaying turbulence. Most notably, it determines the velocity correlation functions and the energy spectrum [Migdal(2024h)]. The computations are rather heavy; in addition to above distribution of the radius $R$ of the regular star polygons, they involve some QFT technology, such as path integrals in the WKB limit, related to the classical trajectory for the Ising spin density $\alpha(k/N)$ and the functional determinant for harmonic fluctuations around this trajectory, computed by the zeta regularization method.
The results for vorticity correlation functions in Fourier space are Mellin-Barnes integrals, with a meromorphic function involving the exponential of another integral arising from the zeta-regularized log-determinant of the aforementioned quadratic form of harmonic fluctuations around the classical path. These integrals, though, are calculable with high precision by Mathematica ® .
These two key observables—the energy spectrum in Fourier space and the velocity moments in coordinate space—are not independent. They are two different representations of the same underlying velocity correlation function. The energy spectrum $E(k,t)$ is directly related to the Fourier transform of this function, while the second-order velocity moment $\langle(\Delta\boldsymbol{v})^{2}\rangle(r,t)$ is its manifestation in real space. Therefore, the theoretical prediction for the scaling exponents is unified, governing both the shape of the energy spectrum in $k$-space and the scaling of the velocity moments in $r$-space.
The second moment of the velocity difference, for instance, is given by an infinite series of power laws, determined by a Mellin-Barnes integral:
$\langle(\Delta\boldsymbol{v})^{2}\rangle(r,t)\propto\frac{1}{t}\sum_{\text{poles }p_{i}}\text{Res}[V(p_{i})]\left(\frac{r}{\sqrt{t}}\right)^{p_{i}}$ (46)
The function $V(p)$, derived from the theory (see Appendix C), has poles that define the universal intermittency exponents, $p_{i}$. This spectrum of indices is remarkable: it contains not only rational numbers but also an infinite series of complex-conjugate pairs, $p_{n}=7\pm it_{n}$, where $\frac{1}{2}\pm it_{n}$ are the nontrivial zeros of the Riemann zeta function. This prediction [Migdal(2024h)], which is in excellent agreement with high-resolution direct numerical simulations (DNS), establishes a deep and unexpected connection between turbulence and number theory.
A direct test of this predicted functional form is the effective index, $f(x)=\langle r\partial_{r}\log(\Delta v^{2})\rangle$, where $x=r/L(t)$. This test is shown in Figure 2 (Top Right), which compares the parameter-free theoretical curve (orange line) to the latest 4K DNS data (blue line) [Sreenivasan and Rodhiya, Akash(2026)]. The agreement is excellent across the primary range.
Deviations are visible at large $\log(x)$, which are expected. This right-hand part of the plot, where $x=r/L(t)$ is large, is contaminated by two distinct systematic effects. First, at any given time, the scales $r$ are approaching the finite simulation box size, leading to lattice artifacts. Second, the time-averaging includes data from early simulation times, when $L(t)$ is still small and influenced by the initial conditions; this small $L(t)$ also shifts the data to large $x$.
The true statistical errors are understood to be much smaller than the displayed error bands, which are inflated by these systematic, non-statistical artifacts. We display the full curve, including these deviations, to avoid any temptation to discard data that does not fit the theory. Instead, these deviations at large $x$ are themselves understood as known, physical consequences of the simulation’s constraints. Even with these effects, the theoretical curve remains within approximately two standard errors of the data, indicating strong agreement.
Furthermore, the log-log slope of $E(t)$ versus $L(t)$ in the DNS data shows a slight, systematic deviation from the leading-order theoretical exponent of 5/2-5/2. This deviation is not a flaw; it is perfectly explained by the theory.
As shown in Figure 2 (Bottom), the leading-order term (gray dash-dotted line) acts as an asymptote, while the complete theoretical curve (orange line)—which includes all sub-leading exponents from the numerically evaluated Mellin-Barnes integral—provides a precise match to the $Re_{\lambda}=145$ DNS data over the entire range. This fit, which matches thousands of data points using only two physical scale parameters (for overall energy and length), provides a high level of confidence in the solution. It visually confirms that the observed energy decay law is composed of the "ground state" term ($L_{M}^{-5/2}$) plus a hierarchy of calculable, quantum-like corrections, just as the theory predicts.
Presumably, the minor deviations from the theoretical curve at the far left end ($\log L_{M}<3.2,\log E>2$) are explained by the early stages of energy decay, when statistical equilibrium has not yet been achieved. The initial K41 spectrum would have contaminated the data, leading to deviations from the turbulent attractor.
4.2 Key Prediction 2: The Geometry of Mixing
The Euler ensemble also dictates the behavior of a passive scalar (e.g., temperature or a dye) advected by the turbulent flow [Migdal(2025e)]. For a localized initial condition, the scalar does not spread into a simple Gaussian cloud. We derived the loop equation for the evolution of the passive scalar $T(r,t)$ from conventional diffusion-advection equations, coupled with the loop equations for the velocity circulation. The resulting solution for the scalar density in the extreme turbulent limit is quite unexpected. Instead of a smooth Gaussian cloud, it forms a series of expanding, quantized concentric shells (see Figure 3. The radial profile of the scalar concentration is piecewise-parabolic, and the shell structure is organized by the Euler totient function, which counts the number of valid star polygons of a given complexity. This sharp, geometric prediction offers a clear target for future experiments.
Figure 3: Log–log plot of the universal function $\xi^{2}\Phi\left(\left\lfloor\frac{1}{\xi}\right\rfloor\right)$ where $\xi=\frac{2\pi r}{\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)}$. Here $\Phi(n)=\sum_{0<p<n}\varphi(p)$ is the Euler totient summatory function.
5 Application II: Magnetohydrodynamic (MHD) Turbulence
The loop space solution can be generalized to more complex systems, such as magnetohydrodynamics (MHD), which describes the dynamics of conducting fluids, such as astrophysical plasmas [Migdal(2025d)]. The system is now described by two coupled loop functionals, one for the velocity circulation ($\Gamma_{v}$) and one for the magnetic vector potential circulation ($\Gamma_{a}$), each governed by an interacting Euler ensemble.
The solution depends critically on the magnetic Prandtl number, $\Pr=\nu/\eta$, the ratio of kinetic viscosity to magnetic resistivity. The theory predicts a first-order phase transition at $\Pr=1$. (see Figure 4)
  • For $\Pr<1$ (e.g., liquid metals), the turbulence is dominated by hydrodynamic fluctuations.
  • For $\Pr>1$ (e.g., hot plasmas), the system exhibits two branches: a stable regime where magnetic and kinetic energies are balanced, and a metastable regime characterized by growing magnetic fluctuations.
This sharp, falsifiable prediction provides new insight into the behavior of astrophysical and laboratory plasmas. The scale $|f(\Pr)|$ of the wave vector in the spectrum is the following function of the Prandtl number
$|f(\Pr)|=\begin{cases}\sqrt{1+3\Pr}&\text{ if }\Pr<1\\ 1+\Pr-\sqrt{\Pr^{2}-\Pr}&\text{ if }\Pr>1\end{cases}$ (47)
Figure 4: Three phases of the wavevector scale $|f(\Pr)|$ in MHD decaying turbulence. The red-dashed line represents a metastable phase.
6 The Yang-Mills Connection and Universal Duality
The final step reveals the framework’s profound universality. The Yang-Mills gradient flow, a central tool for understanding the vacuum structure and confinement problem in QCD, is governed by a non-Abelian version of the same loop dynamics [Migdal(2025c)]. Applying our framework, we find exact solutions for the evolution of the Wilson loop.
In the Yang–Mills gradient flow, the loop-space calculus leads not only to the diffusion equation for the Wilson loop but also to its stationary fixed point, which defines the confining string. The analytic solution is provided by the Hodge-dual matrix surface, obtained by the harmonic-map method. This surface satisfies the self-duality relation
$\Sigma_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\alpha\beta}\Sigma_{\alpha\beta},$ (48)
and minimizes the dual area functional under this constraint. Its area $S[C]$ exactly solves the non-perturbative fixed-point Yang–Mills loop equation,
$\partial_{\mu}\frac{\delta S[C]}{\delta\sigma_{\mu\nu}(\theta)}=0,$ (49)
ensuring that $\exp\left(-\kappa S[C]\right)$ is annihilated by the loop operator. This construction yields a parity-even, factorized form of the Wilson functional,
$W[C]=\exp\left(-\kappa(S_{+}[C]+S_{-}[C])\right),$ (50)
where $S_{\pm}[C]$ are the areas of the analytic self-dual and anti-self-dual surfaces bounded by the same loop.
For planar loops, the dual area is $|S|=2\sqrt{2}|D|$, reproducing the confining area law, while for general nonplanar or twisted loops, the shape dependence distinguishes it from the Euclidean minimal surface. The harmonic-map solution [Migdal(2025a)] clarifies the geometry: the Hodge-dual surface corresponds to a harmonic embedding $X:\mathcal{D}\!\to\!M_{N}(\mathbb{C})$ with the induced conformal metric $g_{ab}=e^{\varphi}\delta_{ab}$, where the Liouville field $\varphi$ governs local scale fluctuations. This analytic result provides the first explicit geometric realization of the Yang–Mills confining string and establishes a bridge between gauge theory, dual string geometry, and the loop-space formulation of turbulence.
The fact that the very same quantized number-theoretic structure governs the universal attractor in both classical fluid turbulence and fundamental gauge theory is the strongest evidence of a deep unifying principle. This duality—between a strongly coupled, chaotic field theory and a weakly coupled, solvable string theory on a discrete target space—is a new manifestation of a principle familiar from the AdS/CFT correspondence.
7 Suggestions for Future Experiments and Tests of the Theory
The loop space calculus framework provides a rich set of new, quantitative, and falsifiable predictions that can be addressed by the next generation of experiments and high-resolution direct numerical simulations (DNS). We outline two primary directions for future work that could provide definitive tests of the theory.
7.1 Visualizing Quantized Scalar Shells
Figure 5: The ’ramp-cliff’ structure in the time trace of temperature fluctuations ($\Delta\theta$) in a heated turbulent jet. These asymmetric patterns, characterized by a gradual rise (the ’ramp’) followed by a sharp drop (the ’cliff’), are a key signature of large-scale coherent structures imprinting on small-scale scalar fields. This empirical observation is analogous to the sawtooth profile predicted for the quantized scalar shells in the Euler ensemble (Figure 3) if you reflect this picture across the right-hand side.( Figure adapted from [Sreenivasan(2019b)]
As detailed in Section 3.3, our theory predicts that a passive scalar released from a localized source in decaying isotropic turbulence will not simply diffuse into a Gaussian cloud. Instead, it should form a series of discrete, expanding concentric shells, whose spacing and density profile are determined by the Euler totient summatory function. The time-domain signature of these shells at a fixed sensor would be a characteristic "sawtooth" decay (see Figure 3).
While existing DNS data has been invaluable for testing spectral properties, most studies have focused on fluctuation statistics rather than the coherent evolution of the mean field. A definitive test of this unique spatial prediction requires a dedicated numerical experiment. To this end, a collaboration has been initiated to solve the passive-scalar advection-diffusion equation concurrently with an existing high-resolution DNS of decaying isotropic turbulence. This project will provide the first direct test for the existence of these quantized shell structures.
Presumably, these shell structures manifest as the "ramp-cliff" structures with similar sawtooth profiles, observed long ago but never completely understood from the microscopic theory ( see [Sreenivasan(2019b)], (see Figure 5)
Figure 6: Bottom: Experimental data from the Max Planck Institute group showing distinct oscillations in the effective index of the second-order velocity structure function at large values of the scaling variable. Top: The theoretical prediction from the Euler ensemble, showing qualitatively similar oscillations. The period of these oscillations is a fundamental prediction of the theory, linked to the imaginary parts of the zeros of the Riemann zeta function.
7.2 Confirming Quantum Interference Oscillations
Perhaps the most surprising prediction of this theory is the emergence of "quantum" interference effects within a purely classical system. The spectrum of scaling exponents derived from the Euler ensemble contains an infinite series of complex-conjugate pairs, tied to the nontrivial zeros of the Riemann zeta function. A complex exponent of the form $p=a+ib$ contributes to correlation functions with terms that oscillate in logarithmic space, i.e., $\sim r^{a}\cos(b\log r)$.
Standard multifractal phenomenology has no mechanism to produce such oscillations. However, remarkably, precisely these types of oscillations have been independently observed in recent, high-precision experiments [Küchler et al.(2023)Küchler, Bewley, and Bodenschatz] on decaying turbulence by the Max Planck Institute group (see Figure 6, Bottom). This data was not published in that paper, but the authors provided it at my request. I shared it with their permission. here.
Figure 7: Left: The turbulent kinetic energy decaying with time in the $4K$ grid simulation with initial Gaussian velocity field distributed according to the K41 spectrum $k^{-5/3}$. After a few hundred time steps, the statistical equilibrium was reached, and energy started decaying in agreement with our $t^{-5/4}$ law. There are other parameters of that stage of decay, with distribution matching our theory, such as the length scale $L(t)\sim\sqrt{t}$ and effective index of the second moment of velocity field which is a nontrivial function of scaling variable $\log(r/\sqrt{t})$ (see Fig. 2). Right: Typical setup of the grid turbulence experiment. The flow enters from the left, passes through the oscillating grid, and creates vorticity. These vortexes interact, exchange energy, and eventually reach statistical equilibrium, with energy constantly dissipated in the bulk while being resupplied by the flow through the grid. In the still Galilean frame, this is a steady process, with energy coming from the left boundary and dissipated in the bulk. The local kinetic energy decays with the distance $z$ from the grid as $z^{-1.25}$, but stays constant with time. In the frame moving with the mean flow velocity, this is decaying turbulence, with energy decaying with time as $(vt)^{-1.25}$.
Our theory predicts a qualitatively similar oscillatory pattern (Figure 6, Top). Crucially, the period of the oscillations observed in the experimental data is of the same order of magnitude as the period predicted by the imaginary part of the first nontrivial zero of the Riemann zeta function ($b\approx 14.13$).
This tantalizing agreement presents a clear and critical target for future high-precision experiments. A definitive confirmation of this oscillatory pattern and a quantitative match with the predicted period would constitute the first experimental evidence of macroscopic "quantum" effects in classical turbulence, directly validating loop space quantum mechanics and its deep connection to number theory.
These quantum effects have nothing to do with quantum turbulence in liquid helium—they follow from the exact mathematical equivalence of our loop equation to the wave equation in loop space, with a non-Hermitian Hamiltonian and viscosity in place of Planck’s constant.
8 A New Perspective on Classical Models in Decaying Turbulence
Grid Turbulence and DNS
A classic setting is grid turbulence (Fig. 7). The landmark experiments of [Comte-Bellot and Corrsin(1966)] measured a decay exponent $1.25\pm 0.01$, close to $5/4$ and inconsistent with the Kolmogorov–Saffman value $6/5$. More recently, DNS on $4\mathrm{K}$ grids with periodic boundaries and K41-type initial spectra (no forcing) were performed by Sreenivasan and collaborators. The simulations relaxed to a decaying state in which the energy decays as $t^{-5/4}$, the integral scale obeys $L(t)\sim t^{1/2}$, and the effective index of the second moment follows the predicted nonlinear curve (Fig. 7,Left), 2).
Relation to Classical Turbulence Paradigms
The loop equations as a route to decaying turbulence represent a departure from customary practice, and they naturally meet resistance from the community. The objection is often framed as follows: rather than resolving the myriad vortical structures and their interactions in physical space, the loop approach reduces the problem to the solution of a stochastic process on regular star polygons (Fig. 1). Can such a one-dimensional description be truly equivalent to three–dimensional vortex dynamics?
A related concern is that any putative solution must address the familiar themes of the classical picture: energy cascade, vortical hierarchies, multifractal scaling, and intermittency. Where do these ideas sit in the loop description? Our position is simple. Some widely used assumptions and scaling templates do not apply to the decaying turbulence regime; others are modified or generalized in the loop picture. Below, we explain how the main questions are answered within the theory.
The Physical Interpretation of Loop Space Geometry
The polygons in Figure 1 are not meant to be seen in a flow visualization; they are a convenient basis in loop space. Their role is analogous to an unobservable but predictive degrees of freedom in other areas of physics (such as quark confinement): they lead to quantitative, testable predictions, such as the decay of turbulent kinetic energy as $t^{-\frac{5}{4}}$.
Duality and the Recasting of Complexity
The key is duality: the same physics can admit complementary mathematical descriptions. In one representation, the dynamics look intricate (waves and eddies); in another, they are encoded by a simpler stochastic object. Just as strong coupling in one theory can map to weak coupling in a dual theory, the turbulent limit of the velocity dynamics maps to a WKB-like limit of the Euler ensemble, which is exactly solvable at infinite Reynolds number. In this sense, the “complexity” is not lost; it is recast.
The complexity is thus not diminished, but rather transformed into a different mathematical language. It shifts from the statistical phenomenology of multifractal exponents and moment-closure hierarchies to the deep arithmetic structures of number theory, revealing a connection between turbulence and concepts such as Euler totients and the Riemann zeta function.
The Nature of the String Theory Duality
A natural question arises from the perspective of string theory: In what sense is the Euler ensemble a "string theory" if the loop $C$ is not the boundary of a worldsheet?
The connection is one of duality, analogous to the relationship between position and momentum space in quantum field theory. The Euler ensemble solution for the loop functional, Eq. (32), is precisely the generating functional of a 1D quantum field theory—a string theory—living on the loop itself.
Let us break down the analogy:
  • The String Field: The momentum loop, $f_{\mu}(\theta)$, which executes a random walk on a discrete target space (the vertices of star polygons), plays the role of the string’s position field.
  • The Source/Momentum: The physical loop, $\dot{C}_{\mu}(\theta)$, does not represent the string itself. Instead, it acts as a source or an external momentum field that couples to the string’s position, $f_{\mu}(\theta)$.
  • The Amplitude: The loop functional $\Psi[C]$ is, therefore, the string’s vacuum-to-vacuum amplitude in the presence of this external momentum source. Averaging over the Euler ensemble is the path integral over all allowed string configurations.
This correspondence manifests a momentum-space duality. The loop $C$ lives in our physical space, but it acts as a momentum-space probe for a string residing in a separate, dual space.
This interpretation elegantly resolves the apparent paradox of a discrete spectrum. The quantization and discreteness (the star polygons, the number-theoretic structure) exist in the dual momentum space of the string. The physical loops $C$ in our world remain continuous and smooth. The theory does not assume nor predict discrete loops in physical space, but rather a discrete spectrum of statistical correlations that are probed by continuous loops.
Thus, the Euler ensemble is not a conventional string theory where the loop is a boundary. It is a dual string theory with a discrete target space, where the physical loop $C$ functions as the momentum variable in its generating functional.
The Hodge–Dual Surface as a Theoretical Area Law in QCD
Why introduce a new area law when lattice QCD already confirms the conventional one? The answer lies in distinguishing a numerical observation from an analytic solution. Two arguments are central.
  • Analytic foundation. Numerical simulations can test or falsify models but cannot derive the law itself from first principles. The analytic Hodge–dual matrix surface, constructed by the harmonic–map method, exactly solves the fixed–point Yang–Mills loop equation [Migdal(2025a)]. It provides the theoretical explanation for the confining area law rather than an empirical fit, offering a genuine analytic counterpart to the numerical area law of lattice QCD.
  • Beyond current lattice reach. Present lattice resolutions are insufficient to test the distinctive, shape–dependent predictions of the Hodge–dual surface. Existing calculations mostly involve planar Wilson loops, for which the dual area is $2\sqrt{2}$ times the Euclidean area—a factor that can be absorbed into the string tension. The crucial test would involve nonplanar, “twisted” contours where the dual and Euclidean minimal surfaces yield different areas. Such computations remain demanding but would provide a definitive check of the analytic theory once feasible.
In this sense, the Hodge–dual matrix surface does not replace the lattice area law but explains its origin and predicts measurable deviations for nonplanar geometries, establishing an analytic framework for the QCD confining string. Being an exact analytic solution, it also provides a natural starting point for quantizing the string, beginning from this holomorphic minimal-surface solution. The resulting quantized theory offers a systematic path to computing the hadron spectrum from first principles, extending far beyond the phenomenology of the linearly rising potential.
Kolmogorov Scaling Violation in the Decaying Regime
With respect to the decaying regime, the classical K41 template is not supported by data at the level often assumed. Kolmogorov’s original constant–dissipation ansatz was a model assumption, and even he (with Obukhov) introduced fluctuations (log–normal) soon after. Modern experiments and DNS [Yakhot and Zakharov(1993), Yakhot and Sreenivasan(2004), Sreenivasan and Yakhot(2021)] show significant, systematic deviations from K41 (and from its log–normal variant). In decaying turbulence, experiments [Panickacheril John et al.(2022a)Panickacheril John, Donzis, and Sreenivasan] deviate even more strongly, despite attempts to fit or trim the data; for example, the energy spectrum departs from $k^{-\frac{5}{3}}$ over many decades, and recent measurements [Küchler et al.(2023)Küchler, Bewley, and Bodenschatz] report clear departures of the "effective index" of $\langle\Delta v^{2}\rangle(r)$ from the K41 value of $2/3$.
Multifractals and Conformal Symmetry
Motivated by analogies with critical phenomena, multifractal models posit power laws with anomalous exponents in an inertial range and can organize deviations from K41 in forced turbulence [Parisi and Frisch(1985), Yakhot and Sreenivasan(2004), Sreenivasan and Yakhot(2021)]. In the incompressible turbulence, however, there is no theoretical basis for conformal symmetry: the conservation constraints of a CFT would assign equal dimensions d1=2d-1=2 to divergence-less fields $\boldsymbol{v},\boldsymbol{\omega}$. This relation is incompatible with $\boldsymbol{\omega}=\nabla\times\mathbf{v}$, which would lead to dimensions differing by $1$ rather than being equal. More importantly, the decaying spectra are not straight lines on log–log plots; they bend universally (cf. Fig. 2 (top) of [Küchler et al.(2023)Küchler, Bewley, and Bodenschatz]), indicating that single power laws—whether K41 or multifractal—are insufficient. The loop theory predicts nonlinear, universal functions of $\log k\sqrt{t}$ across several decades, consistent with the observed curvature.
A Generalization of Multifractal Laws
While not conformally invariant, loop space theory provides a microscopic basis for and a generalization of the multifractal paradigm: the multifractal models introduce a continuous spectrum of anomalous exponents phenomenologically, but the loop theory derives, from first principles, an infinite, discrete, and arithmetic spectrum of intermittency exponents. The Mellin transform of the vorticity correlation is found to be meromorphic, yielding the set of scaling exponents shown below:
indexes of velocity correlation
$\begin{array}[]{c|c}\text{Index}&\text{Condition}\\ \hline\cr-1&\\ 0&\\ \hline\cr 2n&n\in\mathbb{Z},n\geq 1\\ 5/2&\\ 11/2&\\ \frac{15+4n}{2}&n\in\mathbb{Z},n\geq 0\\ 7\pm\imath t_{n}&n\in\mathbb{Z}\end{array}$ (59)
where $\frac{1}{2}\pm\imath t_{n}$ are the nontrivial zeros of the Riemann zeta function.
This result fundamentally refines the multifractal idea. The leading exponents in this spectrum, consistent with numerical data, can be viewed as the "ground state" of the scaling law, describing the dominant behavior. The infinite series of subsequent exponents then dictates a hierarchy of calculable, quantum-like corrections. Recent DNS [Sreenivasan(2023)] report slopes consistent with these predictions, in contrast to the classical scaling templates of Kolmogorov or Saffman.
The 4/5 Law, Spectral Flux, and Vorticity
A central pillar of the classical cascade narrative is the Kolmogorov 4/5 law. In its familiar scalar form, the law relates the third-order moment of the longitudinal velocity increment to the separation distance $r$, i.e., $\langle(\delta v_{L})^{3}\rangle=-\frac{4}{5}\epsilon r$. This linear dependence on $r$ is widely interpreted as direct evidence for a constant, scale-independent energy flux through the inertial range.
A more complete analysis, however, requires examining the law’s full tensorial structure, $S_{3}^{\alpha\beta\gamma}(\mathbf{r})=\langle v_{\alpha}(\mathbf{0})v_{\beta}(\mathbf{0})(v_{\gamma}(\mathbf{r})-v_{\gamma}(\mathbf{0}))\rangle$. The exact solution to the Karman-Howarth equation shows that this tensor is strictly linear in the separation vector $\mathbf{r}$ rather than its length $|\mathbf{r}|$.
This linearity has a direct and crucial consequence in Fourier space. Its Fourier transform is proportional to the gradient of a delta function, $\nabla\delta(\mathbf{k})$, meaning its entire physical content is localized to the largest scale of the system ($\mathbf{k}=0$). The law, therefore, provides a rigorous constraint on large-scale dynamics but does not describe a process of energy transfer between scales.
Furthermore, this tensor is fundamentally irrotational. The law is therefore silent on the statistics of vorticity, which is the essential ingredient of the turbulent cascade. Indeed, the most direct third-order mixed correlation involving vorticity, $\langle vv\omega\rangle$, vanishes identically in isotropic turbulence for purely kinematical reasons of parity invariance. The 4/5 law, thus, offers no constraint on the non-trivial vorticity correlations that would be required to describe a scale-to-scale energy transfer, nor does it limit the scaling properties of vorticity correlations.
The Vanishing Flux and the Dissipative Anomaly
The present framework calls into question not only the K41 scaling laws but also the conventional concept of a "Kolmogorov energy flux." The rapid decay of the energy spectrum predicted by the theory, in accordance with bounds established by Sulem and Frisch [Sulem and Frisch(1975)], corresponds to a vanishing Kolmogorov flux in the turbulent limit. The apparent paradox of finite dissipation in the absence of a spectral flux is resolved by the dissipative anomaly, a mechanism not accounted for in that earlier work.
This mechanism suggests an alternative physical picture to the Richardson cascade. Rather than a scale-by-scale transfer, energy injected at the largest scales (e.g., from boundaries) is dissipated directly on fine-scale, viscous micro-structures. The exact expressions for dissipation and the energy spectrum derived from the loop space approach [Migdal(2024h)] are consistent with the dissipative anomaly and this picture of direct dissipation, but not with a constant-flux energy cascade.
Heisenberg’s Dissipative Range Model
Heisenberg’s $k^{-7}$ proposal for a dissipative subrange, taken up by Chandrasekhar, was a model of its time [Heisenberg(1948), Chandrasekhar(1949)]. As Sreenivasan recounts [Sreenivasan(2019a)], von Neumann already noted in 1949 the lack of experimental support. In our Mellin spectrum, there is no pole between 13/2-13/2 and $8\pm\mathrm{i}t_{n}$, and the "effective index" approaches 7/2-7/2 rather than 7-7. In other words, the nontrivial dynamics persist across what used to be separated as “inertial” and “dissipative” ranges. The resulting universal decaying spectrum spans several decades, and the "effective index" of $\langle\Delta v^{2}\rangle(r)$ varies smoothly from $2$ down to 0, with no plateau at $2/3$. These predictions match DNS and experiments across the entire turbulent range without adjustable dimensionless parameters.
The Question of Universality and Relevance to Steady Flows
The applicability of results from forced simulations to universal turbulence characteristics requires careful consideration. With finite forcing kept active, the resulting statistical state can depend significantly on the specifics of the forcing scheme (e.g., its spatial structure, correlation time, and the mechanism of energy injection), potentially violating universality (see, e.g., the discussion in [Ishihara et al.(2009)Ishihara, Gotoh, and Kaneda]). In particular, large-wavelength forcing, especially if targeting the potential (irrotational) velocity component, can potentially seed non-universal large–scale structures influenced by long–range correlations $\langle vv\rangle(r)\sim r^{\alpha}\to\infty$. In such cases, the simulated flow may reflect properties of the driving mechanism more than the intrinsic fluid dynamics, akin to altering the taste of soup with the greasy spoon used to stir it.
By contrast, the loop equation isolates the vorticity sector, whose correlations inherently decay with distance and are expected to be less sensitive to boundary conditions and to details of distant forcing. This separation helps identify a universal decaying–turbulence attractor. Crucially, this attractor is directly relevant even to ostensibly steady turbulent flows, such as Feynman’s pipe flow, experimental grid turbulence, and jets. These flows typically exhibit spatial decay downstream from the generation source. By Galilean invariance, in a frame moving with the mean flow velocity $v$, this spatial decay is equivalent to the temporal decay problem addressed by our theory, with downstream distance $z$ corresponding to time $t=z/v$ and the boundary condition at the energy source, such as the inlet of the tube, becoming an initial condition at $t=0$ when a given element of water passes the energy source. The observed spatial energy decay in grid turbulence, $E(z)\sim z^{-5/4}$ [Comte-Bellot and Corrsin(1966)], provides a compelling example, matching the temporal $E(t)\sim t^{-5/4}$ decay predicted by the Euler ensemble. Thus, the study of temporally decaying turbulence offers a first-principles approach to understanding the universal aspects of these physically ubiquitous, spatially developing flows, addressing the long-standing challenges highlighted by Feynman.
Ergodicity on the Attractor
In Newtonian mechanics, the evolution of the probability distribution in phase space (the Liouville equation $\partial_{t}\rho=\{H,\rho\}$) is solved by any function of the Hamiltonian $H$. Physical arguments select the Gibbs distribution, $\rho\propto\exp\left(-\beta H\right)$. This argument relies on the properties of the stationary solution itself, not directly on the ergodic hypothesis—the (still unproven) conjecture that a trajectory of a conservative Hamiltonian system eventually covers its energy surface uniformly. Despite the lack of a general proof for ergodicity, the Gibbs distribution has been successfully applied for over a century.
In our theory, the loop diffusion equation replaces the Liouville equation, and the fixed-point (decaying) solution that replaces the Gibbs distribution is the Euler ensemble. The analog of the ergodic hypothesis is the conjecture that every distinct state of the Euler ensemble contributes equally to long-time averages, implying that each state is visited with equal frequency over time. We have not proven this conjecture; the closest approach involved obtaining a transcendental equation for the spectrum of the time decay of deviations from the Euler ensemble [Migdal(2023)]. A proof of ergodicity (or convergence to the attractor) would need to demonstrate the stability of the Euler ensemble, i.e., the positivity of the real parts of all the decay indices in its spectrum.
Lacking such a proof, we turn to physical and numerical experiments. These consistently show convergence from initial states with large vorticity towards a solution whose statistical properties closely match the predictions derived from the Euler ensemble, including decay indices and spectra, as discussed in previous sections [Comte-Bellot and Corrsin(1966), Panickacheril John et al.(2022b)Panickacheril John, Donzis, and Sreenivasan, Sreenivasan and Rodhiya, Akash(2026)]. While the stability and ergodic properties of the Euler ensemble remain a mathematical challenge, this empirical agreement provides strong support for its role as the relevant invariant measure for decaying turbulence.
Summary.
In the decaying turbulence, single–power scaling laws and cascade arguments do not describe the observed behavior. The loop equation replaces these templates with a microscopic, parameter–free description: universal nonlinear spectral shapes, an arithmetic spectrum of indices, and a diffusion in loop space that organizes the statistics. The evidence from classic experiments and modern DNS supports this picture across the turbulent range.
9 Conclusion and Outlook
The loop space calculus offers a new perspective on a class of strongly nonlinear problems in theoretical physics. The approach is based on reformulating the dynamics from local fields to non-local loop observables, recasting the governing nonlinear equations into a universal linear diffusion equation in the space of loops. This transformation enables direct analytical treatment by applying the functional Fourier transform, leading to a solvable algebraic equation.
For decaying turbulence, this has led to a parameter-free solution—the Euler ensemble—which is shown to be dual to a solvable string theory. A central result is the discovery of a deep connection between the flow’s spatial statistics and its temporal evolution. Where phenomenological models treat them separately, this framework derives both the spectrum of spatial intermittency exponents and the spectrum of temporal decay exponents from a single, underlying number-theoretic structure related to the zeros of the Riemann zeta function. The leading exponents in these spectra correspond to observed scaling and decay laws, while the infinite series of subsequent exponents provides a hierarchy of calculable, quantum-like corrections. The theory’s predictions are consistent with available numerical data and offer clear targets for future experimental tests.
The appearance of the same mathematical solutions in the context of the Yang-Mills gradient flow suggests a broad applicability of these methods. These applicability points towards a common structure underlying the statistical behavior of these disparate physical systems. Promising directions for future research include applying this approach to problems in compressible turbulence, cosmology, and the loop dynamics of Einstein gravity.
Among these avenues, the theory’s formulation of the confining string is of particular interest for quantum chromodynamics. The framework yields an exact solution to the QCD loop equations for non-intersecting loops in the form of a Hodge-dual minimal surface [Migdal(2025c)]. This geometric object, by its very construction, satisfies the loop equations in four-dimensional space, unlike the conventional Euclidean minimal surface, which violates the full set of Yang-Mills loop equations.
The analytic Hodge–dual matrix surface derived by the harmonic–map method [Migdal(2025a)] provides the geometric core of the Yang–Mills fixed–point solution, completing the dual correspondence between turbulence and confinement. Beyond this analytic construction, the holomorphic reduction yields an explicit Confinement Theorem. Define the spectral ratio
$r[C]$ $=\frac{\lambda_{3}+\lambda_{4}}{\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4}}\in\Big[\tfrac{1}{2},\,1\Big],$ (60)
where $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\lambda_{4}$ are the ordered eigenvalues of the Douglas–Gram tensor
$G_{\mu\nu}[C]$ $=\sum_{n>0}n\,\hat{C}_{n,\mu}\,\hat{C}_{n,\nu},\qquad\hat{C}_{n,\mu}=\tfrac{1}{2\pi}\!\int_{0}^{2\pi}\!C_{\mu}(\theta)e^{-in\theta}\,d\theta.$ (61)
From the analytic solution, the Hodge–dual minimal area of the surface and the Euclidean minimal area of the same boundary loop are expressed as
$S_{\chi}[C]$ $=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$ $S_{\mathrm{E}}[C]$ $=\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4},$ (62)
so that their ratio directly yields the spectral ratio $r[C]$. Hence the effective string tension,
$\sigma_{\mathrm{eff}}[C]$ $=\frac{S_{\chi}[C]}{S_{\mathrm{E}}[C]}=2\sqrt{2}\,r[C],$ (63)
obeys the universal bounds
$\sqrt{2}\ \leq\ \sigma_{\mathrm{eff}}[C]\ \leq\ 2\sqrt{2},$ (64)
for all loops. The numerical value of the lower bound is simple, but its existence—independent of loop size and shape—is highly nontrivial and constitutes the confinement statement in this framework. All residual shape dependence of the string tension is encoded by the spectral ratio $r[C]$: planar loops (rank–2 $G$) saturate the upper bound $r[C]=1$, while isotropic spectra saturate the lower bound $r[C]=\tfrac{1}{2}$. This explicit formula for the shape dependence of the quark string tension opens the way to computing the rising Regge trajectories in QCD from first principles.
While the same mathematical formalism connects turbulence, gauge theory, and string duality, these correspondences should not be interpreted as a unification of the underlying physical systems but rather as a shared geometric language linking universal features of their non-linear dynamics.
Sometimes, understanding familiar physical phenomena requires learning new mathematical methods, as happened historically with planetary motion and chemical reactions. In my pursuit of a solution to turbulence, I found it necessary, over decades of exploration, to learn relevant aspects of geometry and number theory. This process eventually led to a geometric understanding of loop-space diffusion and the parallel transport operator: the loop-space calculus. This new perspective then naturally revealed the turbulent attractor rooted in number theory—the Euler ensemble. Assuming that this solution withstands further scrutiny and validation, it invites a shared effort to explore this rich mathematical landscape, echoing the spirit of Euler, who brought together distinct fields of mathematics to describe natural laws.
Acknowledgements
I am grateful for insightful discussions regarding this theory with Herman Verlinde, who offered several inspiring remarks. Conversations with Katepalli Sreenivasan concerning the physics of turbulence, as observed in direct numerical simulations and laboratory experiments, were invaluable in clarifying the role and significance of loop space diffusion.
Supported by the Simons Foundation (2019–2024).
Appendix A Essentials of the loop space calculus
The loop space calculus is a variational framework that operates entirely within the manifold of smooth loops. It replaces singular operations (e.g., adding infinitesimal loops) with coincident limits of smooth "dot derivatives" $\frac{\delta}{\delta\dot{C}(\theta)}$. This appendix summarizes the key definitions, identities, and results of the calculus.
The full derivations and technical details can be found in our recent paper [Migdal(2025b)].
The Wilson loop and non-abelian Stokes theorem
The area derivative is defined by the non-abelian Stokes theorem. For a detailed derivation, see [Migdal(2025b)], Sec. 3.3.
$\delta W[C]=\oint d\theta\dot{C}_{\mu}(\theta)\delta C_{\nu}(\theta)\frac{\delta W[C]}{\delta\sigma_{\mu\nu}(\theta)}$ (65)
A.1 Operator representation of the Wilson loop
The loop functional can be represented as a trace of a path-ordered exponential of covariant derivative operator. This subsection proves the fundamental operator identity that enables this. The full derivation, originally presented in [Migdal(2025b)] (Sec. 3.4), is reproduced here with all necessary mathematical details. We also fix some typos in that paper.
Theorem 1 (Operator-Holonomy Identity).
Let $D_{\mu}(x)=\mathbb{I}_{G}\partial_{\mu}+A_{\mu}(x)$ be the covariant derivative at an arbitrary base point $x\in\mathbb{R}^{d}$, where $\partial_{\mu}$ is the partial derivative in Hilbert space and $A_{\mu}(x)$ is a matrix-valued operator in the group space $G$, and $\mathbb{I}_{G}$ is a unity matrix in that group space. Let $C$ be a smooth, closed loop parameterized by $\theta\in[0,2\pi]$ with $\dot{C}_{\mu}(\theta)$ as its velocity, such that $C(\theta)=x+\int_{0}^{\theta}d\theta^{\prime}\dot{C}_{\mu}(\theta^{\prime})$, and $\int_{0}^{2\pi}d\theta^{\prime}\dot{C}_{\mu}(\theta^{\prime})=0$.
The path-ordered exponential of the covariant derivative operator $D_{\mu}(x)$ along this loop is equal to the Wilson loop (holonomy) along the path $C(\theta)$ multiplied by the identity operator $\mathbb{I}$ in the Hilbert space.
$\mathbb{P}\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)D_{\mu}(x)\right)=\mathbb{P}\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)A_{\mu}(C(\theta))\right)\otimes\mathbb{I}$ (66)
Proof.
The proof relies on discretizing the path and using the "disentangling" identity, which is a consequence of Feynman’s operator calculus [Feynman(1951)].
1. Discretization
The path-ordered exponential on the L.H.S. is formally defined as the limit of a product integral:
$\mathbb{P}\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)D_{\mu}(x)\right)=\lim_{N\to\infty}\prod_{k=N\to 1}\exp\left(\Delta\theta_{k}\dot{C}_{\mu}(\theta_{k})D_{\mu}(x)\right)$ (67)
where $\Delta\theta_{k}\to 0$ and the product is ordered from right to left (i.e., $k=1$ is the rightmost operator). Let $dC_{\mu}(\theta)=d\theta\dot{C}_{\mu}(\theta)$.
2. The Infinitesimal Disentangling
We analyze a single infinitesimal factor $\exp\left(dC_{\mu}D_{\mu}(x)\right)$. Using the Lie-Trotter product formula, $e^{A+B}=e^{A}e^{B}+O([A,B])$, we can split the operator:
$\exp\left(dC_{\mu}(\partial_{\mu}+A_{\mu}(x))\right)=\exp\left(dC_{\mu}\partial_{\mu}\right)\exp\left(dC_{\mu}A_{\mu}(x)\right)+O(d\theta^{2})$ (68)
The $O(d\theta^{2})$ terms vanish in the $N\to\infty$ limit. The $\exp\left(dC_{\mu}\partial_{\mu}\right)$ operator is an infinitesimal translation operator, $T_{dC}$. We now use the fundamental operator identity that defines a translation:
$T_{dC}\,f(x)=f(x+dC)\,T_{dC}$ (69)
Applying this to our $A_{\mu}(x)$ operator, we find the "hopping" identity:
$\exp\left(dC_{\mu}\partial_{\mu}\right)\exp\left(dC_{\mu}A_{\mu}(x)\right)=\exp\left(dC_{\mu}A_{\mu}(x+dC)\right)\exp\left(dC_{\mu}\partial_{\mu}\right)$ (70)
This is the "disentangling" from (A7) in the text. It allows us to move the (Abelian) derivative operator to the right, at the cost of shifting the argument of the (non-Abelian) $A_{\mu}$ operator.
3. Iteration and Re-ordering
We now apply this identity iteratively to the full product. Let $C_{k}=C(\theta_{k})=x+\int_{0}^{\theta_{k}}d\theta^{\prime}\dot{C}(\theta^{\prime})$. By periodicity, $C_{N}=x$.
$\prod_{k=N\to 1}\exp\left(dC_{k}\cdot D(x)\right)\approx\prod_{k=N\to 1}\left(\exp\left(dC_{k}\cdot A(C_{k})\right)\exp\left(dC_{k}\cdot\partial_{x}\right)\right)$ (71)
$=\left(e^{dC_{N}\cdot A(C_{N})}e^{dC_{N}\cdot\partial_{x}}\right)\left(e^{dC_{N-1}\cdot A(C_{N})}e^{dC_{N-1}\cdot\partial_{x}}\right)\dots\left(e^{dC_{1}\cdot A(C_{N})}e^{dC_{1}\cdot\partial_{x}}\right)$ (72)
We repeatedly "hop" all the $\exp\left(dC_{k}\cdot\partial_{x}\right)$ terms to the far right. Each time a $\partial_{x}$ operator passes an $A(C_{N})$ operator, it shifts its argument. The crucial point is that all $\partial_{x}$ operators are Abelian and commute with each other, while all $A(C_{k})$ operators are non-Abelian and remain path-ordered. Furthermore, the resulting argument $x_{k}$ of each $A$ in the product at kthk-th place will be
$x_{k}=C_{N}+\sum_{n=N\to k}dC_{n}=C_{k}$ (73)
This process separates the product into two distinct, ordered parts:
$=\left(\prod_{k=N\to 1}e^{dC_{k}\cdot A(C_{k})}\right)\left(\prod_{k=N\to 1}e^{dC_{k}\cdot\partial_{x}}\right)$ (74)
Taking the $N\to\infty$ limit, this becomes:
$=\left(\mathbb{P}\exp\left(\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)A_{\mu}(C(\theta))\right)\right)\left(\exp\left(\int_{0}^{2\pi}dC_{\mu}\partial_{\mu}\right)\right)$ (75)
4. The Closed Loop Condition
The first term is, by definition, the Wilson loop $W[C]$ (the holonomy), which is a c-number matrix. The second term is the identity operator in Hilbert space, because the loop $C$ is closed:
$\int_{0}^{2\pi}dC_{\mu}=\int_{0}^{2\pi}d\theta\dot{C}_{\mu}(\theta)=C_{\mu}(2\pi)-C_{\mu}(0)=0$ (76)
Therefore, the operator becomes:
$\exp\left(\left(\int_{0}^{2\pi}dC_{\mu}(\theta)\right)\partial_{x_{\mu}}\right)=\exp\left(0\cdot\partial_{x_{\mu}}\right)=\mathbb{I}$ (77)
This completes the proof. We find that the Hilbert space operator on the L.H.S. is equal to the group space matrix (the Wilson loop) multiplied by the identity operator $\mathbb{I}$. ∎
The area derivative
The field strength $F_{\mu\nu}$ is generated by the antisymmetric part of the second dot derivative. For details, see [Migdal(2025b)], Sec. 3.6, Eq. (25).
$\frac{\delta W(C(.),\tau)}{\delta\sigma_{\mu\nu}(t)}=\frac{\delta}{\delta\dot{C}_{\left[\mu\right.}(t-)}\frac{\delta}{\delta\dot{C}_{\left.\nu\right]}(t+)}W(C(.),\tau)=$
$\left\langle\frac{1}{N}\hbox{tr}\,F_{\mu\nu}(C(t))P\exp\left(\mathrm{i}\int_{t}^{t+2\pi}ds\dot{C}_{\mu}(\theta)D_{\mu}(C(t))\right)\right\rangle$ (78)
The covariant derivative
The covariant derivative of the field strength is generated by the discontinuity of the dot derivative, $\partial_{\mu}(t)\equiv\frac{\delta}{\delta\dot{C}_{\mu}(t-)}-\frac{\delta}{\delta\dot{C}_{\mu}(t+)}$. For details, see [Migdal(2025b)], Sec. 3.7, Eq. (31).
$\partial_{\mu}(t)\frac{\delta W[C]}{\delta\sigma_{\mu\nu}(t)}=\left\langle\hbox{tr}\,\left[D_{\mu},F_{\mu\nu}(C(t))\right]\right$.
$\left.P\exp\left(\mathrm{i}\int_{t}^{t+2\pi}d\theta\dot{C}_{\mu}(\theta)D_{\mu}(C(t))\right)\right\rangle$ (79)
Annihilation property
The dot derivative discontinuity $\partial_{\mu}(t)$ annihilates the Wilson loop. For details, see [Migdal(2025b)], Sec. 3.9, Eq. (33).
$\partial_{\mu}(t)W(C(.),\tau)=0;$ (80)
The Leibniz rules
The area derivative acts as a first-order differential operator, satisfying the standard Leibniz rules. For details, see [Migdal(2025b)], Sec. 3.10, Eqs. (39-41).
$\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}(A[C]B[C])$ $=\frac{\delta A[C]}{\delta\sigma_{\mu\nu}(\theta)}B[C]+A[C]\frac{\delta B[C]}{\delta\sigma_{\mu\nu}(\theta)};$ (81)
$\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}F(A[C])$ $=F^{\prime}(A[C])\frac{\delta A[C]}{\delta\sigma_{\mu\nu}(\theta)};$ (82)
$\partial_{\mu}\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}(A[C]B[C])$ $=0,\quad\text{if A, B solve the fixed-point eq.}$ (83)
The Bianchi identity
The calculus contains a kinematical Bianchi identity. For details, see [Migdal(2025b)], Sec. 3.11, Eq. (43).
$e_{\alpha\mu\nu\lambda}\left(\frac{\delta}{\delta\dot{C}_{\alpha}(\theta+)}-\frac{\delta}{\delta\dot{C}_{\alpha}(\theta-)}\right)\frac{\delta}{\delta\dot{C}_{\left[\nu\right.}(\theta-)}\frac{\delta}{\delta\dot{C}_{\left.\mu\right]}(\theta+)}=0;$ (84)
The nonsingular loop equation
The dot derivatives combine to form the loop operator $\mathcal{L}_{C}$, yielding a closed, linear diffusion equation for the Wilson loop functional. For details, see [Migdal(2025b)], Sec. 3.12, Eqs. (45-48).
$\partial_{\tau}W[C,\tau]$ $=\mathcal{L}_{C}W[C,\tau]$ (85)
$\mathcal{L}_{C}$ $=\oint d\theta\dot{C}_{\nu}(\theta)\hat{L}_{\nu}(\theta);$ (86)
$\hat{L}_{\nu}(\theta)$ $=T_{\nu}^{\alpha\beta\gamma}\frac{\delta^{3}}{\delta\dot{C}_{\alpha}(\theta-0)\delta\dot{C}_{\beta}(\theta)\delta\dot{C}_{\gamma}(\theta+0)};$ (87)
$T_{\nu}^{\alpha\beta\gamma}$ $=\delta_{\alpha\beta}\delta_{\gamma\nu}+\delta_{\gamma\beta}\delta_{\alpha\nu}-2\delta_{\alpha\gamma}\delta_{\beta\nu};$ (88)
Appendix B Back to the static loop equation
We now map the Euler-ensemble solution of the liquid loop equation back to the static loop equation by prescribing a loop motion in momentum space that reproduces the missing $\mathbf{v}\times\boldsymbol{\omega}\,$ term. Consider
$\Psi[C,t]=\Big\langle\exp\left(\mathrm{i}\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\mathbf{P}(\theta)\right)\Big\rangle_{\mathcal{E}},$ (89)
$\partial_{t}\mathbf{C}(\theta)=-\,\mathbf{v}_{P}(\theta),$ (90)
$\frac{D}{Dt}\Psi[C,t]=\partial_{t}\Psi[C,t]$
$-\Big\langle\mathrm{i}\oint d\theta\ \partial_{\theta}\mathbf{v}_{P}(\theta)\cdot\mathbf{P}(\theta)\,\exp\left(\mathrm{i}\oint\mathbf{C}^{\prime}\cdot\mathbf{P}\right)\Big\rangle_{\mathcal{E}}.$ (91)
Integrating by parts and using $\mathrm{i}\,\partial_{\theta}\mathbf{P}\Rightarrow-\,\delta/\delta\mathbf{C}$ on the exponential yields the desired static advection operator. The velocity is expressed in terms of vorticity by inverting the curl, given zero divergence.
$\vec{v}=\frac{-1}{\vec{\nabla}^{2}}\vec{\nabla}\times\vec{\omega}$ (92)
With the dictionary (valid inside the momentum loop Anzatz)
$\nabla\;\Rightarrow\;\frac{\partial}{\partial\mathbf{C}(\theta)}=\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}-\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)},$ (93)
$\hat{\omega}(\theta)\;\Rightarrow\;-\,\mathrm{i}\,\nu\,\frac{\delta}{\delta\boldsymbol{\sigma}(\theta)}$
$=-\,\mathrm{i}\,\nu\,\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}\times\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}.$ (94)
we find
$\nabla\Rightarrow\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}-\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}\Rightarrow-\,\mathrm{i}\,\Delta\mathbf{P},$ (95)
$\boldsymbol{\omega}\Rightarrow\mathrm{i}\nu\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}\times\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}$
$\Rightarrow\mathrm{i}\nu\mathbf{P}_{\mathrm{mid}}\times\Delta\mathbf{P},$ (96)
so that
$\mathbf{v}_{P}=\nu\frac{\Delta\mathbf{P}\times(\mathbf{P}_{\mathrm{mid}}\times\Delta\mathbf{P})}{(\Delta\mathbf{P})^{2}}\quad\stackrel{{\scriptstyle\mathcal{E}}}{{\Longrightarrow}}\quad\mathbf{v}_{P}=\nu\mathbf{P}_{\mathrm{mid}}.$ (97)
The extra term in momentum space is then
$\mathrm{i}\oint d\theta\ \partial_{\theta}\mathbf{v}_{P}\cdot\mathbf{P}=-\,\mathrm{i}\oint d\theta\ \mathbf{v}_{P}\cdot\partial_{\theta}\mathbf{P}$
$=\mathrm{i}\nu\oint d\theta\ \mathbf{P}_{\mathrm{mid}}\cdot\partial_{\theta}\mathbf{P},$ (98)
which vanishes: the smooth part $\mathbf{P}_{\mathrm{mid}}$ of $\mathbf{P}$ gives an integral of total derivative $\frac{1}{2}\partial_{\theta}\mathbf{P}_{\mathrm{mid}}^{2}$ over the closed loop, and the jump contributions $\mathbf{P}_{\mathrm{mid}}\cdot\Delta\mathbf{P}$ vahish locally by the Euler-ensemble constraints. Equivalently, solving for the moving momentum loop
$\mathbf{C}(\theta,t)=\mathbf{C}_{0}(\theta)+\big(\sqrt{2\nu(t+t_{0})}-\sqrt{2\nu t_{0}}\big)\,\mathbf{F}_{\mathrm{mid}}(\theta)$ (99)
one finds that the time-dependent term cancels as in the liquid equation. Thus the The Euler ensemble solves both the liquid and static loop equations.
Appendix C Velocity Correlations and the Spectrum of Exponents
The Euler ensemble solution yields an explicit formula for the second moment of the velocity difference as a Mellin transform:
$\langle(\Delta\boldsymbol{v})^{2}\rangle(r,t)=\frac{\tilde{\nu}^{2}}{\nu t}\int_{\epsilon-i\infty}^{\epsilon+i\infty}\frac{dp}{2\pi i}V(p)\left(\frac{r}{\sqrt{\tilde{\nu}t}}\right)^{p}$
where $\tilde{\nu}$ is a turbulent viscosity and the kernel $V(p)$ is a meromorphic function given by:
$V(p)=-\frac{f(-1-p)\zeta(\frac{13}{2}-p)\csc(\frac{\pi p}{2})}{16\pi^{2}(p+1)(2p-15)(2p-5)\zeta(\frac{15}{2}-p)}$
Here, $\zeta(s)$ is the Riemann zeta function and $f(z)$ is a calculable entire function defined in Appendix K in [Migdal(2024h)]. This function involves other functions $A(\Delta),B(\Delta),C(\Delta)$ that are defined in previous Appendices to that paper. These functions were precomputed in Mathematica ® and stored as fifth order interpolation through the dense table of results of high-precision numerical integration using Mathematica ® .
The poles of $V(p)$ give the spectrum of intermittency exponents, which includes both rational values and complex values determined by the zeros of $\zeta(s)$. The entire function f(1p)f(-1-p) drops from this spectrum of poles, contributing only to the residues. This is why the spectrum reduces to rational numbers and the zeros of the zeta function, – the basic objects of number theory.
Complete details of analytical and numerical computations involved in this large project are summarized in collection of Mathematica ® notebooks [Migdal(2024a), Migdal(2024b), Migdal(2024c), Migdal(2024d), Migdal(2024e), Migdal(2024f), Migdal(2024g)], which can be freely downloaded to verify these computations.
Appendix D The Hodge-Dual Matrix Surface
This section summarizes the definition of the Hodge-dual matrix-valued minimal surface. The full derivations and properties are detailed in Sec. 4.2 of [Migdal(2025b)].
Parametric Definition
The surface is embedded in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ with matrix-valued coordinates $X_{\mu}^{A}(\xi_{1},\xi_{2})$. For details, see [Migdal(2025b)], Sec. 4.2, Eq. (82).
$X_{\mu}=\tau_{A}X^{A}_{\mu}(\xi_{1},\xi_{2})=X^{0}_{\mu}(\xi_{1},\xi_{2})+\tau_{a}X^{a}_{\mu}(\xi_{1},\xi_{2})$ (100)
Scalar Area and Hodge-Duality
The area functional is defined with a Hodge-duality constraint. For details, see [Migdal(2025b)], Sec. 4.2, Eqs. (83-86).
$d\Sigma_{\mu\nu}$ $=d\xi_{1}d\xi_{2}\Sigma_{\mu\nu};$ (101)
$\Sigma_{\mu\nu}$ $=e_{lm}\partial_{l}X^{A}_{\mu}\partial_{m}X^{A}_{\nu};$ (102)
$|S|$ $=\int_{\mathcal{D}}\sqrt{\frac{1}{2}(\Sigma_{\mu\nu})^{2}};$ (103)
$\Sigma_{\mu\nu}$ $=\frac{1}{2}e_{\mu\nu\alpha\beta}\Sigma_{\alpha\beta};$ (104)
Dirichlet Boundary Conditions
The surface boundary is mapped from the loop $C(\theta)$ using ’t Hooft projectors. For details, see [Migdal(2025b)], Sec. 4.2, Eqs. (91-97).
$X^{A}_{\mu}(1,\theta)=\Lambda^{A}_{\mu\nu}C_{\mu}(\theta);$ (105)
$\Lambda^{0}_{\mu\nu}=\delta_{\mu\nu};\quad\Lambda^{a}_{\mu\nu}=\bar{\eta}^{a}_{\mu\nu};$ (106)
$\eta^{i}_{\mu\nu}=(\delta_{4\nu}\delta_{i\mu}-\delta_{4\mu}\delta_{i\nu})+e_{i\mu\nu 4};$ (107)
$\bar{\eta}^{i}_{\mu\nu}=(\delta_{4\nu}\delta_{i\mu}-\delta_{4\mu}\delta_{i\nu})-e_{i\mu\nu 4};$ (108)
$\frac{1}{2}e_{\mu\nu\lambda\rho}\eta^{i}_{\lambda\rho}=\eta^{i}_{\mu\nu};\quad\frac{1}{2}e_{\mu\nu\lambda\rho}\bar{\eta}^{i}_{\lambda\rho}=-\bar{\eta}^{i}_{\mu\nu};$ (109)
Conformal Gauge Condition
The conformal gauge is fixed by a condition on the induced metric at the boundary. For details, see [Migdal(2025b)], Sec. 4.2, Eq. (107).
$g_{lm}|_{\partial D}=\left(\partial_{l}\bar{X}^{A}_{\mu}\,\partial_{m}X^{A}_{\mu}\right)_{\partial D}\propto\delta_{lm}.$ (110)
Appendix E The loop equation for the minimal surface and its solution
This section summarizes how the Hodge-dual surface provides an exact, non-singular solution to the fixed-point loop equation. The full derivation is in Sec. 4.3 of [Migdal(2025b)].
Variational Principle
The area functional is defined as the minimum of a constrained action. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (111-113).
$|S|$ $=\min_{\lambda,X}\int d^{2}\xi\left(\sqrt{\frac{1}{2}\hbox{tr}\,\Sigma\cdot\Sigma^{T}}+\lambda_{i}\bar{\eta}^{i}_{\mu\nu}\Sigma_{\mu\nu}\right);$ (111)
$t_{\mu\nu}(r,\theta)$ $=\frac{\Sigma_{\mu\nu}}{\sqrt{2\hbox{tr}\,\Sigma\cdot\Sigma^{T}}}$ (112)
Area Derivative
The variation of the area functional $\delta|S|$ yields the area derivative. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (121, 122).
$\delta|S|$ $=\int d\theta\delta C_{\alpha}(\theta)\dot{C}_{\beta}(\theta)2T_{\alpha\beta}(1,\theta);$ (113)
$\frac{\delta|S|}{\delta\sigma_{\alpha\beta}(\theta)}$ $=2T_{\alpha\beta}(1,\theta)=2(t_{\alpha\beta}+\lambda_{i}\bar{\eta}^{i}_{\alpha\beta});$ (114)
The Hodge-Duality Constraint
The Hodge-duality of the area derivative requires the Lagrange multipliers to vanish at the boundary. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (123, 124).
$\bar{\eta}^{a}_{\mu\nu}\,T_{\mu\nu}(1,\theta)$ $=0\quad\forall a=1,2,3;$ (115)
$\lambda_{i}(1,\theta)$ $=0;$ (116)
The Solution of the Loop Equation
The area derivative of the Hodge-dual surface is, by construction, proportional to the self-dual tensor $t_{\mu\nu}$. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (125).
$\frac{\delta}{\delta\dot{C}_{\left[\mu\right.}(\theta-)}\frac{\delta}{\delta\dot{C}_{\left.\nu\right]}(\theta+)}|S|=2t_{\mu\nu}(1,\theta);$ (117)
This identity, combined with the kinematical Bianchi identity (from Appendix A), proves that the loop operator annihilates the area functional. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (126).
$\partial_{\mu}\frac{\delta|S|}{\delta\sigma_{\mu\nu}(\theta)}\propto e_{\mu\nu\lambda\rho}\left(\partial_{\alpha}(\theta)\right)\frac{\delta}{\delta\dot{C}_{\lambda}(\theta-)}\frac{\delta}{\delta\dot{C}_{\rho}(\theta+)}|S|\equiv 0$ (118)
Therefore, the exponential of the Hodge-dual area is an exact, non-singular solution to the fixed-point loop equation. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (127).
$\mathcal{L}_{\nu}(t)\,\exp\left(-\kappa S[C]\right)=\partial_{\mu}\left(\frac{\delta S[C]}{\delta\sigma_{\mu\nu}}\exp\left(-\kappa S[C]\right)\right)=0$ (119)
Planar Loop minimal surface
The Hodge-dual minimal surface can be analytically found for an arbitrary non-intersecting planar loop. The main result is the following theorem. We present this theorem here, correcting the proof in [Migdal(2025b)], where there was a typo in the boundary conditions for $Y$.
Theorem 2.
The minimal dual area for any non-intersecting planar loop equals $2\sqrt{2}$ times the planar area inside this loop.
Proof.
Let us use the conformal invariance of the dual area functional (101) and map the unit circle onto the given planar bounding loop $\mathbb{S}_{1}\mapsto C$. After that, the minimal surface can use the coordinates $\xi=(x_{1},x_{2})$ inside that planar loop $C$ as internal coordinates of the surface. In the context of our surface embedded into $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$, this parametrization reduces to
$X^{A}_{l}=\Lambda^{A}_{lm}\,\xi_{m},\quad l,m=1,2,$ (120)
$X^{A}_{\dot{l}}=Y^{A}_{\dot{l}}(\xi),\quad\dot{l}=3,4,$ (121)
$\xi\in\mathcal{D},\quad C=\partial\mathcal{D},$ (122)
where $\Lambda^{A}_{\mu\nu}$ are defined in (105), and the new field $Y^{A}_{\dot{l}}(\xi)$ satisfies the boundary conditions:
[…]
The area functional is
$|S|=\int_{\mathcal{D}}d^{2}\xi\sqrt{\tfrac{1}{2}\Sigma_{lm}^{2}+\Sigma_{l\dot{l}}^{2}+\tfrac{1}{2}\Sigma_{\dot{l}\dot{m}}^{2}},$ (124)
$\Sigma_{lm}=e_{ab}\Lambda^{A}_{al}\Lambda^{A}_{bm},$ (125)
$\Sigma_{l\dot{l}}=e_{ab}\Lambda^{A}_{al}\,\partial_{b}Y^{A}_{\dot{l}},$ (126)
$\Sigma_{\dot{l}\dot{m}}=e_{ab}\,\partial_{a}Y^{A}_{\dot{l}}\,\partial_{b}Y^{A}_{\dot{m}}.$ (127)
With the Dirichlet boundary values (123) the constrained (self–dual) minimizer of the area functional (124) is the linear map
$X^{A}_{\mu}(\xi)=\Lambda^{A}_{\mu m}\,\xi_{m},\qquad\xi\in D.$ (128)
This embedding automatically satisfies both the Hodge–duality constraint and the corrected boundary conditions. Its area tensor is constant,
$\Sigma_{\mu\nu}=\sum_{A}\Lambda^{A}E(\Lambda^{A})^{\!T}=2\!\begin{pmatrix}0&1&0&0\\[-3.0pt] -1&0&0&0\\[-3.0pt] 0&0&0&1\\[-3.0pt] 0&0&-1&0\end{pmatrix}$, (129)
which is self–dual ($\Sigma=\Sigma$) and yields $\Sigma_{l\dot{l}}=0$. The Euler–Lagrange system is identically satisfied for linear $X^{A}_{\mu}(\xi)$. The corresponding area density is $\sqrt{\tfrac{1}{2}\Sigma_{\mu\nu}\Sigma_{\mu\nu}}=2\sqrt{2},$ so that
$|S|=\!\!\int_{D}\!d^{2}\xi\,2\sqrt{2}=2\sqrt{2}\,|D|.$ (130)
Hence the minimal dual area for any non–intersecting planar loop equals $2\sqrt{2}$ times the Euclidean area enclosed by the loop, and the linear map (128) provides the unique regular self–dual minimizer consistent with the boundary data. ∎
Area Inequalities for Confinement
For an arbitrary loop $C$, the self-dual surface area $|S|$ is bounded by the unconstrained minimal area in the embedding space. These bounds are crucial for confinement. The full proofs are provided in [Migdal(2025b)] (Sec. 4.8).
Lower Bound
Theorem 3.
The area of the constrained matrix surface (i.e., with the duality constraint) is not smaller than that of the minimal unconstrained matrix surface with the same bounding curve. This follows from the fact that the constrained configuration space is a subset of the unconstrained space. See [Migdal(2025b)], Sec. 4.8, Eq. (168).
$\inf_{X\in\mathbb{R}^{4}\otimes\mathbb{R}^{4},\,\Sigma=\pm*\Sigma}|S|\geq\inf_{X\in\mathbb{R}^{4}\otimes\mathbb{R}^{4}}|S|.$ (131)
Upper Bound
Theorem 4.
The area of the unconstrained matrix surface is positive but not larger than four times the minimal surface area in Euclidean space ($A_{4}[C]$) with the same bounding curve $C\subset\mathbb{R}^{4}$. The proof relies on the inequality $\left|\Sigma_{\mu\nu}\right|\leq\sum_{A}\sqrt{\det\lVert g^{A}_{lm}\rVert}$. This leads to the final bound: See [Migdal(2025b)], Sec. 4.8, Eq. (172).
$0<\min_{X\in\mathbb{R}^{4}\times\mathbb{R}^{4}}|S|\leq 4A_{4}[C].$ (132)
For a planar loop, the Hodge-dual area is $2\sqrt{2}A_{4}[C]\approx 2.83A_{4}[C]$, which respects this bound.
Parity and the Final Solution
A key conceptual issue, first pointed out by E. Witten, is that a single Hodge-dual surface with a fixed chirality ($\lambda=\pm 1$) would violate parity. The full analysis is in [Migdal(2025b)] (Sec. 4.9).
To restore parity and preserve the crucial factorization property for separated loops, the solution must be the exponential of the sum of both the self-dual and anti-self-dual areas. See [Migdal(2025b)], Sec. 4.9, Eq. (176).
$W[C]=\exp\left(-\kappa\,(\,|S^{+}[C]|+|S^{-}[C]|\,)\right).$ (133)
This form is parity-even and solves the fixed-point loop equation because the loop operator, $\partial_{\mu}\frac{\delta}{\delta\sigma_{\mu\nu}(\theta)}$, annihilates each term in the exponent independently via the Bianchi identity (from Appendix A).
Conclusion. This provides a parity-preserving, factorizable, and exact analytic solution for the fixed point of the Yang-Mills gradient flow.